Grade 12Algebra

In a college, out of 4320 students, 2300 are girls. Find the ratio of
Number of girls to the total number of students.

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

Given
Total number of students = 4320
Number of girls = 2300
Number of boys = 4320–2300
= 2020
Ratio of number of girls to the total number of students = 2300 / 4320
= (2 × 2 × 5 × 115) / (2 × 2 × 5 × 216)
= 115 / 216

Thanks